OpenStudy (anonymous):

what is the volume of an oblique cone with 9in. radius and a height of 12in.

OpenStudy (anonymous):

The volume of a cone is 1/3(Area of Base)(height) = 1/3 π r2 h so, 1/3 pi * 81 * 12 = 324π in^3

OpenStudy (anonymous):

thank you thatmathguy

OpenStudy (anonymous):

np